# and Threats in Multimodal Medical RAG Systems

### v1

@@ Introduction
Medical Vision-Language Models (Med-VLMs), which integrate visual and textual modalities, are becoming increasingly central to numerous clinical tasks, including radiology report generation, Visual Question Answering (VQA), and diagnostic decision support (Yuan et al. 2023; Hartsock and Rasool 2024; Chen et al. 2024a). By jointly encoding medical images (e.g., X-rays, MRIs) and clinical text (e.g., questions, reports), these models offer a unified framework for multimodal understanding in healthcare.
Despite their promise, the trustworthiness of Med-VLMs remains a critical concern. Hallucinations, confident but clinically incorrect outputs, pose a serious threat in medical applications (Ji et al. 2023), where even a minor factual error in generated content can lead to severe or even life-threatening consequences. Improving factual consistency and grounding remains an open challenge in clinical AI systems (Miura et al. 2023).
To address this issue, RAG has emerged as a widely adopted solution. By retrieving relevant external knowledge, such as image-report pairs or medical literature (Jain and et al. 2021), RAG augments the generation process with domain-specific evidence, effectively reducing hallucinations and improving factual correctness (Zhao et al. 2024; Wu et al. 2024a). Recent work, such as (Xia et al. 2024), has demonstrated that RAG significantly boosts both the accuracy and reliability of medical large vision-language models when properly adapted to the domain.
However, while RAG reduces hallucinations, it also introduces a new and underexplored vulnerability: sensitivity to data poisoning. In real-world deployments, external knowledge bases often undergo periodic updates, including user contributions or automated data pipelines (Jafarzadeh and Ensan 2025). This semi-open structure opens the door to adversaries who can inject harmful, misleading, or semantically conflicting image-text pairs (Carlini and et al. 2023) into the retrieval corpus, without needing access to model weights or prompts.
Ensuring the security of medical AI models is urgent (Finlayson and et al. 2019) - contamination of external medical knowledge can severely compromise their credibility and clinical value. Thus, safeguarding RAG-based systems from retrieval-stage manipulation is not just a technical issue but a clinical imperative.
@@
Medical Multimodal RAG enhances parametric knowledge by retrieving relevant medical texts and clinical images from an external Knowledge Base (KB) (Chen et al. 2024b), aiding more accurate and grounded generation in the biomedical domain. Inspired by recent advances in multimodal RAG (Zhao et al. 2023), we develop a medical multimodal RAG pipeline consisting of a multimodal medical KB, a domain-specific retriever, a reranker, and a generator.
@@
In the context of medical image analysis, diffusion models have to be first adapted (fine-tuned) to the medical domain for generating accurate synthetic radiographs  (Bluethgen et al. 2024; Dutt et al. 2024b) . Diffusion probabilistic models employ a two-phase procedure: a forward diffusion process and a reverse denoising process (Zhang et al. 2025; Yang et al. 2023). In the forward process, a data sample $x_{0}$ drawn from the true data distribution $q(x_{0})$ is gradually corrupted over $T$ time steps by adding Gaussian noise. Each step follows a predefined noise schedule $\beta_{t}$, resulting in the following Markov process:
@@
VA focus on manipulating the image modality. First, we employ a pre-trained diffusion model (Stable Diffusion V 1-4) (Rombach et al. 2022) fine-tuned on the MIMIC dataset (Dutt et al. 2025) to generate clinically accurate synthetic chest X-rays.This model conditions on a medical text prompt describing clinical characteristics (e.g., “no acute cardiopulmonary abnormality”) and synthesizes corresponding X-ray images that conform to those descriptions.
@@
Implementation Details. We use LLaVA-Med1.5 7B (Li, Wong et al. 2023) as the backbone model. In the multimodal RAG framework, we employ CLIP (Radford et al. 2021) and OpenCLIP (Cherti et al. 2023) as retrievers to fetch relevant medical image-text pairs. For re-ranking and generation, Qwen VL Chat (Bai et al. 2023) and LLaVA (Li, Wong et al. 2023) models are utilized to ensure contextual accuracy and robust output generation. The diffusion model for adversarial visual attacks uses full fine-tuning on the MIMIC-III dataset to synthesize realistic but adversarial medical images. Training and fine-tuning are conducted on four NVIDIA A100 GPUs with mixed precision, employing the AdamW optimizer (learning rate 2e-5, batch size 16, for 10 epochs). Hyperparameters for each stage, retrieval, re-ranking, and generation, are validated using a held-out validation set.
@@
Evaluation Datasets. We evaluate on two widely-used medical imaging benchmarks: (1) IU X-ray  (Pavlopoulos, Kougia, and Androutsopoulos 2019), which contains chest X-ray images aligned with radiology reports; and (2) MIMIC-CXR  (Johnson et al. 2019), a large-scale clinical chest X-ray dataset with aligned question-answer pairs. To simulate poisoning attacks, we create synthetic adversarial image-text pairs using diffusion-based image generation and negation-flip techniques, which are then injected into the external knowledge base to test model resilience.
@@ Results
Vulnerability Landscape of AI medicine. AI-driven tools enhance diagnostics by interpreting complex medical data but are also prone to adversarial threats. Medical AI models across domains like radiology and ECG analysis have shown vulnerabilities to both white-box and black-box attacks (e.g., FGSM, PGD) (Kaviani, Han, and Sohn 2022; Puttagunta, Ravi, and Nelson Kennedy Babu 2023). Medical data’s specificity may heighten these risks. Broader concerns include LLM-aided phishing and data breaches compromising patient privacy (Hassanin and Moustafa 2024; Nemec Zlatolas, Welzer, and Lhotska 2024; Dutt 2025). Addressing these threats requires secure, transparent frameworks with ongoing evaluation and fairness (Dutt et al. 2024a), alongside defences like adversarial training (Bountakas et al. 2023).
@@
Text and Image Data Poisoning. Among the diverse attack methodologies, data poisoning emerges as a particularly insidious threat, targeting the integrity of the model’s training data itself (Cong et al. 2024; Chao et al. 2024). This is especially relevant for modern AI, including medical systems, which increasingly rely on large-scale, often web-sourced, multimodal datasets (text and images) (Yang et al. 2024). Recent research demonstrates the growing sophistication of poisoning attacks tailored for multimodal contexts. Frameworks like MM-POISONRAG (Ha et al. 2025) reveal how multimodal RAG systems, which integrate text and images from external knowledge bases, can be compromised by injecting targeted misinformation (LPA) or disruptive irrelevant content (GPA) designed to manipulate the cross-modal retrieval process (Ha et al. 2025). Similarly, stealthy attacks like Shadowcast (Xu et al. 2024) create visually plausible but malicious image-text pairs to manipulate VLM outputs subtly (He et al. 2024), while others target text-to-image diffusion models or induce LLMs to generate harmful content via fragmented data insertion (Wang et al. 2022). The practicality and stealth of these methods highlight that reliance on external or large-scale data introduces significant poisoning risks, necessitating advanced data validation and defense strategies (Wang et al. 2022). This vulnerability is particularly pertinent when considering architectures like RAG, designed specifically to leverage external knowledge.
@@
RAG in Med-LVLMs. Retrieval-Augmented Generation (RAG) improves Med-LVLMs by integrating external medical knowledge (e.g., literature, databases) to reduce hallucinations and enhance output accuracy in tasks like VQA and report generation (Zhao et al. 2024; Yuan et al. 2023; Hartsock and Rasool 2024; Chen et al. 2024a). Systems like MMed-RAG and RULE (Xia et al. 2024) refine domain-specific retrieval and knowledge integration. However, reliance on external sources introduces new attack vectors. Knowledge poisoning, as seen in MM-POISONRAG, can compromise RAG by injecting malicious content into retrievable data (Taofeek et al. 2022). Ensuring secure and validated retrieval is thus essential for trustworthy deployment in healthcare.
@@ Conclusion
Building safe multimodal medical RAG systems means defending against the three specific threats revealed in this study. To curb Textual Attacks—where false captions or reports are injected to mislead retrieval (Wan et al. 2023)—every new text should pass automatic fact-checking against trusted medical ontologies, negation detection, and language-model consistency scoring, with low-confidence items routed to experts. For Visual Attacks, which rely on synthetic yet plausible images to sway the retriever (Bortsova et al. 2021), incoming images are screened with perceptual hashing, diffusion-artifact detectors, and out-of-distribution filters; flagged images are down-ranked or held for human review. To stop Cross-Modal Conflict Injection, where paired images and texts are semantically misaligned to confuse the generator (Bortsova et al. 2021; Wan et al. 2023), the pipeline computes an image–text entailment score and drops pairs that show high contradiction. Across all defences we keep signed provenance logs, enable instant rollback of suspect knowledge-base snapshots, stream real-time telemetry with alert thresholds tied to clinical risk, and empower clinicians to veto any retrieval influencing patient care. These targeted measures, combined with regular red-team drills and strict HIPAA/GDPR compliance, help keep the knowledge base trustworthy and the model’s answers reliable. Crucially, our framework is highly flexible and adaptable, because all safety hooks attach at well-defined API boundaries, the retrieval component can be hot-swapped—e.g., replacing a dense-vector RAG with a drop-in GraphRAG that navigates medical ontologies—without re-engineering the defence stack. This modularity stands in contrast to the recent Medical Graph RAG (Wu et al. 2024b), whose safety guarantees hinge on a fixed graph schema. While their method chiefly addresses hallucination through graph grounding, our framework covers a broader adversarial surface (textual, visual, and cross-modal) and couples detection with operational safeguards such as provenance logging and clinician veto, thereby providing stronger end-to-end safety within comparable latency budgets.

### v2

@@ Introduction
Medical Vision-Language Models (Med-VLMs), which integrate visual and textual modalities, are becoming increasingly central to numerous clinical tasks, including radiology report generation, Visual Question Answering (VQA), and diagnostic decision support (yuan2023ramm; hartsock2024vision; chen2024survey). By jointly encoding medical images (e.g., X-rays, MRIs) and clinical text (e.g., questions, reports), these models offer a unified framework for multimodal understanding in healthcare.
Despite their promise, the trustworthiness of Med-VLMs remains a critical concern. Hallucinations, confident but clinically incorrect outputs, pose a serious threat in medical applications (zhou2023survey), where even a minor factual error in generated content can lead to severe or even life-threatening consequences. Improving factual consistency and grounding remains an open challenge in clinical AI systems (miura2023factual).
To address this issue, RAG has emerged as a widely adopted solution. By retrieving relevant external knowledge, such as image-report pairs or medical literature (jain2021knowledge), RAG augments the generation process with domain-specific evidence, effectively reducing hallucinations and improving factual correctness (zhao2024retrieval; wu2024medical). Recent work, such as (xia-etal-2024-rule), has demonstrated that RAG significantly boosts both the accuracy and reliability of medical large vision-language models when properly adapted to the domain.
However, while RAG reduces hallucinations, it also introduces a new and underexplored vulnerability: sensitivity to data poisoning. In real-world deployments, external knowledge bases often undergo periodic updates, including user contributions or automated data pipelines (jafarzadeh2025evidence). This semi-open structure opens the door to adversaries who can inject harmful, misleading, or semantically conflicting image-text pairs (carlini2023poisoning) into the retrieval corpus, without needing access to model weights or prompts.
Ensuring the security of medical AI models is urgent (finlayson2019adversarial) - contamination of external medical knowledge can severely compromise their credibility and clinical value. Thus, safeguarding RAG-based systems from retrieval-stage manipulation is not just a technical issue but a clinical imperative.
@@
Medical Multimodal RAG enhances parametric knowledge by retrieving relevant medical texts and clinical images from an external Knowledge Base (KB) (chen2024mllmstrongrerankeradvancing), aiding more accurate and grounded generation in the biomedical domain. Inspired by recent advances in multimodal RAG (zhao2023retrieving), we develop a medical multimodal RAG pipeline consisting of a multimodal medical KB, a domain-specific retriever, a reranker, and a generator.
@@
In the context of medical image analysis, diffusion models have to be first adapted (fine-tuned) to the medical domain for generating accurate synthetic radiographs  (bluethgen2024vision; dutt2024parameterefficient) . Diffusion probabilistic models employ a two-phase procedure: a forward diffusion process and a reverse denoising process (zhang2025frdiff; yang2023diffusion). In the forward process, a data sample $x_{0}$ drawn from the true data distribution $q(x_{0})$ is gradually corrupted over $T$ time steps by adding Gaussian noise. Each step follows a predefined noise schedule $\beta_{t}$, resulting in the following Markov process:
@@
VA focus on manipulating the image modality. First, we employ a pre-trained diffusion model (Stable Diffusion V 1-4) (rombach2022high) fine-tuned on the MIMIC dataset (dutt2025memcontrol) to generate clinically accurate synthetic chest X-rays.This model conditions on a medical text prompt describing clinical characteristics (e.g., “no acute cardiopulmonary abnormality”) and synthesizes corresponding X-ray images that conform to those descriptions.
@@
Implementation Details. We use LLaVA-Med1.5 7B (li2023llava) as the backbone model. In the multimodal RAG framework, we employ CLIP (radford2021learning) and OpenCLIP (cherti2023reproducible) as retrievers to fetch relevant medical image-text pairs. For re-ranking and generation, Qwen VL Chat (bai2023qwen) and LLaVA (li2023llava) models are utilized to ensure contextual accuracy and robust output generation. The diffusion model for adversarial visual attacks uses full fine-tuning on the MIMIC-III dataset to synthesize realistic but adversarial medical images. Training and fine-tuning are conducted on four NVIDIA A100 GPUs with mixed precision, employing the AdamW optimizer (learning rate 2e-5, batch size 16, for 10 epochs). Hyperparameters for each stage, retrieval, re-ranking, and generation, are validated using a held-out validation set.
@@
Evaluation Datasets. We evaluate on two widely-used medical imaging benchmarks: (1) IU X-ray  (pavlopoulos2019survey), which contains chest X-ray images aligned with radiology reports; and (2) MIMIC-CXR  (johnson2019mimic), a large-scale clinical chest X-ray dataset with aligned question-answer pairs. To simulate poisoning attacks, we create synthetic adversarial image-text pairs using diffusion-based image generation and negation-flip techniques, which are then injected into the external knowledge base to test model resilience.
@@ Results
Vulnerability Landscape of AI medicine. AI-driven tools enhance diagnostics by interpreting complex medical data but are also prone to adversarial threats. Medical AI models across domains like radiology and ECG analysis have shown vulnerabilities to both white-box and black-box attacks (e.g., FGSM, PGD) (kaviani2022adversarial; puttagunta2023adversarial). Medical data’s specificity may heighten these risks. Broader concerns include LLM-aided phishing and data breaches compromising patient privacy (hassanin2024comprehensive; nemec2024data; dutt2025devil). Addressing these threats requires secure, transparent frameworks with ongoing evaluation and fairness (dutt2024fairtune), alongside defences like adversarial training (bountakas2023defense).
@@
Text and Image Data Poisoning. Among the diverse attack methodologies, data poisoning emerges as a particularly insidious threat, targeting the integrity of the model’s training data itself (cong2024test; chao2024systematic). This is especially relevant for modern AI, including medical systems, which increasingly rely on large-scale, often web-sourced, multimodal datasets (text and images) (yang2024understanding). Recent research demonstrates the growing sophistication of poisoning attacks tailored for multimodal contexts. Frameworks like MM-POISONRAG (ha2025mm) reveal how multimodal RAG systems, which integrate text and images from external knowledge bases, can be compromised by injecting targeted misinformation (LPA) or disruptive irrelevant content (GPA) designed to manipulate the cross-modal retrieval process (ha2025mm). Similarly, stealthy attacks like Shadowcast (xu2024shadowcast) create visually plausible but malicious image-text pairs to manipulate VLM outputs subtly (he2024evilpromptfuzzer), while others target text-to-image diffusion models or induce LLMs to generate harmful content via fragmented data insertion (wang2022threats). The practicality and stealth of these methods highlight that reliance on external or large-scale data introduces significant poisoning risks, necessitating advanced data validation and defense strategies (wang2022threats). This vulnerability is particularly pertinent when considering architectures like RAG, designed specifically to leverage external knowledge.
@@
RAG in Med-LVLMs. Retrieval-Augmented Generation (RAG) improves Med-LVLMs by integrating external medical knowledge (e.g., literature, databases) to reduce hallucinations and enhance output accuracy in tasks like VQA and report generation (zhao2024retrieval; yuan2023ramm; hartsock2024vision; chen2024survey). Systems like MMed-RAG and RULE (xia-etal-2024-rule) refine domain-specific retrieval and knowledge integration. However, reliance on external sources introduces new attack vectors. Knowledge poisoning, as seen in MM-POISONRAG, can compromise RAG by injecting malicious content into retrievable data (taofeek2022cognitive). Ensuring secure and validated retrieval is thus essential for trustworthy deployment in healthcare.
@@ Conclusion
Building safe multimodal medical RAG systems means defending against the three specific threats revealed in this study. To curb Textual Attacks—where false captions or reports are injected to mislead retrieval (turn2file7)—every new text should pass automatic fact‐checking against trusted medical ontologies, negation detection, and language‐model consistency scoring, with low‐confidence items routed to experts. For Visual Attacks, which rely on synthetic yet plausible images to sway the retriever (turn2file4), incoming images are screened with perceptual hashing, diffusion‐artifact detectors, and out‐of‐distribution filters; flagged images are down‐ranked or held for human review. To stop Cross-Modal Conflict Injection, where paired images and texts are semantically misaligned to confuse the generator (turn2file4; turn2file7), the pipeline computes an image–text entailment score and drops pairs that show high contradiction. Across all defences we keep signed provenance logs, enable instant rollback of suspect knowledge‐base snapshots, stream real-time telemetry with alert thresholds tied to clinical risk, and empower clinicians to veto any retrieval influencing patient care. These targeted measures, combined with regular red-team drills and strict HIPAA/GDPR compliance, help keep the knowledge base trustworthy and the model’s answers reliable. Crucially, our framework is highly flexible and adaptable, because all safety hooks attach at well‑defined API boundaries, the retrieval component can be hot‑swapped—e.g., replacing a dense‑vector RAG with a drop‑in GraphRAG that navigates medical ontologies—without re‑engineering the defence stack. This modularity stands in contrast to the recent Medical Graph RAG (wu2024medicalgraphragsafe), whose safety guarantees hinge on a fixed graph schema. While their method chiefly addresses hallucination through graph grounding, our framework covers a broader adversarial surface (textual, visual, and cross‑modal) and couples detection with operational safeguards such as provenance logging and clinician veto, thereby providing stronger end‑to‑end safety within comparable latency budgets.